In a walk a thon to raise money for charity,Elisa walked a certain distance at 5mph and then jog twice that distance at 8mph. Her total time walking and jogging was 2 hours and 15 minutes. How far did Elisa get?
Answer by stanbon(75887)

In a walk a thon to raise money for charity,Elisa walked a certain distance at 5mph and then jog twice that distance at 8mph. Her total time walking and jogging was 2 hours and 15 minutes. How far did Elisa get?
-------
Walk DATA:
rate = 5 mph ; distance = x miles ; time = d/r = x/5 hrs.
------------------------------
Jogging DATA:
rate = 8 mph ; distance = 2x miles ; time = d/r = 2x/8 hrs
-----
Equation:
time + time = 2 1/4 = 9/4 hrs.
x/5 + 2x/8 = 9/4 hrs
---
Multiply thru by 40 to get:
8x + 10x = 90
18x = 90
x = 5 miles
-----
Total Distance = walk + jog = x+2x = 3x = 15 miles
===================
Cheers,
Stan H.
===================
=====================================
Cheers,
Stan